he cash prices of six-month and one-year Treasury bills are 95.0 and 90.0, respectively. A 1.5-year bond that will pay coupons of $5 every six months currently sells for $94.84. Calculate the 1.5-year zero rate (hint, you must compute the 6-month and 1-year zero rates).

a) 13.16%

b) 13.63%

c) 13.81%

d) 13.99%
Answer #1

6 month zero rate=(100/95)^(1/0.5)-1=10.8033%

1 year zero rate=(100/90)^(1/1)-1=11.11%

1.5 year zero rate=(((105)/(94.84-5*0.95-5*0.90))^(1/3)-1)*2=13.99%